OpenStudy (lgbasallote):

A person walks first at constant speed of 5.00 m/s along a straight line from point A to point B and then back along the line from B to A at a constant speed of 3.00 m/s. What is her average speed over the entire trip?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

30/8

OpenStudy (lgbasallote):

15 meters?

OpenStudy (lgbasallote):

why not just 1? i mean it's simpler isnt it?

OpenStudy (shubhamsrg):

total dist = 2d total time = d/5 + d/3 = 8d/15 so avg speed = 2d/(8d/15) so @A.Avinash_Goutham was right!! yet again! :D

OpenStudy (anonymous):

4m/s

OpenStudy (lgbasallote):

so it will be 30/8 seconds?

OpenStudy (anonymous):

When the distance is fixed, just find the harmonic mean.

OpenStudy (lgbasallote):

i mean m/s

OpenStudy (lgbasallote):

what does harmonic mean mean again?

OpenStudy (anonymous):

For two number a and b the harmonic mean is given by 2ab/(a+b)

OpenStudy (kropot72):

Speed = (distance traveled)/(time taken) Total distance traveled = 0 Speed = 0 m/s

OpenStudy (lgbasallote):

where a and b are...?

OpenStudy (lgbasallote):

the given values?

OpenStudy (cwrw238):

speed = distance / time if distance = 1 5 = 1 / 0.2 also 3 = 1 / 0.3333 average speed = 2 / 0.53333

OpenStudy (anonymous):

a=5 and b =3 or vice-versa

OpenStudy (anonymous):

its 3.75 m/s 4 is wrong

OpenStudy (anonymous):

if it's equal " time " u shud use........arithmetic nd equal dist u shud use..........harmonic

OpenStudy (lgbasallote):

this is effective everytime d is equal?

OpenStudy (anonymous):

total distance /total time travelled

OpenStudy (anonymous):

Yes lgba and when time is equal find the arithmetic mean.

OpenStudy (lgbasallote):

arithmetic mean is..?

OpenStudy (anonymous):

2d ----------- d +d -- ---- 5 3 2d --------- 8d ------ 15 2 ---- 8 ------ 15

OpenStudy (anonymous):

2 ---- 8 --- 15 15 ---- 4

OpenStudy (lgbasallote):

lol yeah i got that part =)) btw...a follow-up question is what is the average velocity...the book says zero...is it because displacement is 0?

OpenStudy (anonymous):

yeah it is zero

OpenStudy (shubhamsrg):

yes,,because disp is 0..

OpenStudy (anonymous):

but its speed written in the question???

OpenStudy (lgbasallote):

the first question was average speed

OpenStudy (anonymous):

yeah but there is a difference displacement can come to zero bot distance cannot u should mention adds up whether how many time come and go from same point asdisplacement get zero

OpenStudy (anonymous):

distance adds up*

OpenStudy (anonymous):

lol really? at 5m/s ?

OpenStudy (lgbasallote):

what? o.O

OpenStudy (anonymous):

A person walks first at constant speed of 5.00 m/s <== that guy must be one heck of a power walker haha